OpenStudy (anonymous):

dy/dx = 3 y x^2 and y=3 when x=-2, then what is y?

OpenStudy (help!!!!):

3

OpenStudy (anonymous):

It's a separable equation.

OpenStudy (help!!!!):

can you draw it?

OpenStudy (anonymous):

\[ \begin{array}{rcl} \frac{dy}{dx} &=& 3 y x^2 \\ \frac{1}{y} \frac{dy}{dx} &=& 3 x^2 \\ \int \frac{1}{y} \frac{dy}{dx}dx &=& \int 3 x^2dx \\ u=y&&du = \frac{dy}{dx}dx \\ \int \frac{1}{u} du &=& \int 3 x^2dx \\ \ln(u) +C_2 &=& x^3+C_1 \\ \ln(y) +C_2 &=& x^3+C_1 \\ \ln(y) &=& x^3+C_3 \\ e^{\ln(y)} &=& e^{x^3+C_3} \\ y &=& Ce^{x^3} \\ \end{array} \]Then you have to plug in to solve for \(C\).

OpenStudy (abb0t):

What happen to the 1/3?

OpenStudy (abb0t):

Oh nvm haha

OpenStudy (anonymous):

\[ \int x^2 dx = \frac{x^3}{3}+C \]

OpenStudy (anonymous):

We solve for \(C\):\[ \begin{array}{rcl} y &=& Ce^{x^3} \\ (3) &=& Ce^{(-2)^3} \\ 3 &=& Ce^{-8} \\ 3e^{8} &=& C \\ \end{array} \] Giving us, in the end: \[ y = (3e^8)e^{x^3} = 3e^{x^3+8} \]

OpenStudy (anonymous):

@jenny0828 Get it?

OpenStudy (anonymous):

Very thorough, wio. I appreciate your help!